OpenStudy (anonymous):

prove 8 is a factor of 9^n-1 with mathematical induction

OpenStudy (asnaseer):

do you know the steps involved in mathematical induction?

OpenStudy (anonymous):

yeah i got the first one but i don't know how to do the second

OpenStudy (asnaseer):

so what was your first step?

OpenStudy (anonymous):

prove for n=1

OpenStudy (asnaseer):

yes - and that is simple enough - correct?

OpenStudy (asnaseer):

next step is to ASSUME it is true for some value of n, say for n=k

OpenStudy (asnaseer):

then, using this assumption, you try to prove it is also true for n=k+1

OpenStudy (asnaseer):

understand?

OpenStudy (anonymous):

but i dont have an actual equation

OpenStudy (asnaseer):

if something is a multiple of 8, then how can you represent that mathematically in an equation?

OpenStudy (anonymous):

9^(k+1)-1=8k?

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

/almost/

OpenStudy (anonymous):

8(k+1)

OpenStudy (asnaseer):

first write the original statement as an equation. you are told to prove that \(9^n-1\) is a multiple of 8, i.e you are told to prove:\[9^n-1=8m\]where n and m are integers

OpenStudy (asnaseer):

agreed?

OpenStudy (anonymous):

sure

OpenStudy (asnaseer):

so you showed that for n=1 you have:\[9^1-1=9-1=8\]next you need to ASUME this equation is true for some n=k, i.e. ASSUME it is true to say that:\[9^k-1=8m\]finally, use this result to try to prove that \(9^{k+1}-1\) is also a multiple of 8.

OpenStudy (anonymous):

so 9^(k+1)-1=8m

OpenStudy (asnaseer):

no - the 'm' here is not the same 'm' as in the result for the assumption

OpenStudy (asnaseer):

try to rewrite \(9^{k+1}-1\) in terms of \(9^k-1\) and things should start to make sense

OpenStudy (asnaseer):

HINT: \(9^{k+1}=9\times9^k\)

OpenStudy (anonymous):

i'm lost

OpenStudy (asnaseer):

ok, I'll work through this one for you and hopefully you will see how to proceed with other similar problems

OpenStudy (asnaseer):

\[9^{k+1}-1=9\times9^k-1\]agreed?

OpenStudy (anonymous):

sure

OpenStudy (asnaseer):

therefore:\[9^{k+1}-1=9\times9^k-1=9(9^k-1)+8\]agreed?

OpenStudy (anonymous):

umm sure

OpenStudy (asnaseer):

not sure?

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

how about if I write it as:\[9^{k+1}-1=9\times9^k-1=9\times9^k-9+8=9(9^k-1)+8\]

OpenStudy (anonymous):

its ok keep going

OpenStudy (asnaseer):

so next we use the result from the 2nd step of induction where we ASSUMED that:\[9^k-1=8m\]to get:\[9^{k+1}-1=9\times9^k-1=9(9^k-1)+8=9(8m)+8\]make sense so far?

OpenStudy (anonymous):

u lost me after ''to get"

OpenStudy (asnaseer):

:)

OpenStudy (asnaseer):

ok, let me try again...

OpenStudy (asnaseer):

I'm just going to connect my writing pad to make it easier to explain

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):